OpenStudy (anonymous):

\[\int_c xyds,\] \[C: x=t^2, y=2t, 0 \le t \le 1\] Evaluate the line integral, where C is the given curve.

OpenStudy (experimentx):

\[ ds = \sqrt{ \left ( dx \over dt \right )^2 + \left ( dy \over dt \right )^2} dt\] change everything into t's and integrate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\int_c 2t^3 \sqrt{4t^2 +4} dt\] \[4\int_c t^3 \sqrt{t^2 +1} dt\]

OpenStudy (anonymous):

\[t=sec\theta\] \[dt=sec\theta tan\theta d\theta\] \[4\int_c sec^4 \theta tan^2 \theta d\theta\]

OpenStudy (experimentx):

t = tan \theta

OpenStudy (anonymous):

I sit easier that way?

OpenStudy (anonymous):

*is

OpenStudy (experimentx):

i think integration by parts is also easier u = t^2 v = t sqrt(t^2 + 1)

OpenStudy (anonymous):

Wouldn't trig substitution be easier than the above link?

OpenStudy (experimentx):

oh .. you have to evaluate \int tan^3 x sec ^3 x dx

OpenStudy (experimentx):

and i'm not fond of high power of trigs ...

OpenStudy (anonymous):

\[4\int_c sec^4 \theta tan^2 \theta d\theta\] \[4\int_c (tan^2 \theta +1)tan^2 \theta d\theta\]

OpenStudy (anonymous):

\[4\int_c (tan^4 \theta +tan^2 \theta) d\theta\]

OpenStudy (anonymous):

\[4\int_c tan^4 \theta d \theta +4\int_ctan^2 \theta d\theta\]

OpenStudy (anonymous):

Does this look right so far?

OpenStudy (anonymous):

I think u v substitution is a lot easier. http://www.wolframalpha.com/input/?i=integral+of%28+tangentx%29^4

OpenStudy (anonymous):

integration by parts I mean

OpenStudy (experimentx):

yep ... i would try that.

OpenStudy (anonymous):

I finally got: \[\frac 4 5 (t^2+1)^{\frac 52}-\frac 4 3 (t^2+1)^{\frac 32}\] is the the given curve?

OpenStudy (anonymous):

*this

OpenStudy (anonymous):

yep, so this "c" by the integral is specific for curvature?

OpenStudy (experimentx):

hmm \[ \int_0^1\]

OpenStudy (anonymous):

Thanks!